): Suppose GDP is $1,000 billion, the national debt last year was $500 billion, the interest rate paid on government debt is 7%, and GDP is growing by 5% per year.

              a. If the goal of the government is to hold the debt-GDP ratio constant, what must be the size of               the primary surplus and the size of total budget surplus? Explain your answer.

              b. Suppose the interest rate paid on government debt were to decrease to 5% and the growth rate        of GDP were to increase to 7% while GDP remained at $1,000 billion and the national debt last        year remained at $500 billion, what must be the size of the primary surplus and the size of the               total budget surplus? Explain your answers.
Answer #1

a) Current Debt to GDP ratio
500 / 1000 = 0.5

GDP in period T+1
1000 * 1.05 = 1050

Primary surplus refers to the surplus of the government which excludes interest paid on the government debt.
So the primary budget surplus should be $50 billion.

The total budget surplus is the difference between total income and total expenses of the government. It includes interest paid on government debt.

500 * 0.07 = $35 billion

The total budget surplus should be $15 billion.
50 - 35 = 15

b) Now, the GDP growth rate is 7%
Primary surplus
1000 * 0.07 = $70 billion

Total budget surplus again includes the interest paid on the debt
Interest paid
500 * 0.05 = $25 billion

Total budget surplus
70 - 25 = $45 billion